Question 40 Not yet answered Points out of 1 Sales (75,000 units) Variable costs $750,000 $525,000 $337,500 $262,500 181125 Contribution margin Fixed manufacturing costs Operating income Flag question Eamings before taxe Taxes (at 31% Net Income Shares outstanding 15,000 The Degree of Financial Leverage is Select one: A. 1.29 B. 4.50 C. 3.50 D. 1.32 O
Answer #1



Degree of Financial leverage = Operating income./ Earning before tax = EBIT/EBT = 337,500/262,500 = 1.29

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well